OpenStudy (anonymous):

HEY What is 5(3)n+56= 67+n+y+3-10

OpenStudy (amistre64):

at best its a line....

OpenStudy (dumbcow):

solve for y maybe

OpenStudy (dumbcow):

y = 14n -4

OpenStudy (immanuelv):

the correct answer is : 15 n+56 = n+y+60

OpenStudy (immanuelv):

which can be rewritten as y= 2(7n-2)

OpenStudy (anonymous):

5(3)n+56= 67+n+y+3-10 15n+56=67+n+y-7 15n-n-y=-56-7 14n-y=-63